OpenStudy (anonymous):

If w is the non-real cube root of unity then \[w^{\frac{1}{4}}(w^{\frac{1}{2}}*w^{\frac{1}{6}}*w^{\frac{1}{18}}*w^{\frac{1}{54}})........ infinity\]

hartnn (hartnn):

can u find the sum of geometric sequence 1/2,1/6,1/18..... ??

hartnn (hartnn):

its a/(1-r) = 0.5/(1-1/3)=0.5/(2/3)=3/4 ok?

OpenStudy (anonymous):

i got it....thxxx

hartnn (hartnn):

so w^(1/4)* w^(3/4) = w.

hartnn (hartnn):

welcome :)